OpenStudy (anonymous):

Divide, express in lowest terms: y+2/169 divided by 4-y^2/13

OpenStudy (anonymous):

Seattle sun girl. are you a math teacher....

OpenStudy (anonymous):

\[\left(\begin{matrix}y+2 \\ 169\end{matrix}\right)\left(\begin{matrix}13 \\ 4-y ^{2}\end{matrix}\right)\]

OpenStudy (anonymous):

factorize \[4 - y ^{2} = (2-y)(2+y)\]

OpenStudy (anonymous):

So you can just simplify it from there :)

OpenStudy (anonymous):

You should get \[1/(26-13y)\]